OpenStudy (anonymous):

Find S12 for the series 1 + 2 + 4 + 8 +.. i got 2048

OpenStudy (anonymous):

@Michele_Laino am i right

OpenStudy (michele_laino):

here we have a geometric series, whose first term is a1=1 and constant q= 2

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

so, we can apply this formula: \[\Large {S_{12}} = {a_1}\frac{{1 - {q^{12}}}}{{1 - q}}\]

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

after a substitution, i get this: \[\Large {S_{12}} = {a_1}\frac{{1 - {q^{12}}}}{{1 - q}} = 1 \cdot \frac{{1 - {2^{12}}}}{{1 - 2}} = ...?\]

OpenStudy (anonymous):

-4097

OpenStudy (michele_laino):

I got 4095

OpenStudy (anonymous):

nvm ur right, typo sorry lol

OpenStudy (anonymous):

so its 4095

OpenStudy (michele_laino):

yes!